The number a exceeds the number b by 50 percent. By what percent is the number b smaller than the number a?
Answer by Alan3354(69443)

The number a exceeds the number b by 50 percent. By what percent is the number b smaller than the number a?
--------
a = 1.5b = (3/2)b
b = (2/3)a
--> 1/3 smaller = 33 1/3% smaller